# GraphMend: Code Transformations for Fixing Graph Breaks in PyTorch 2

### v1

@@ Abstract
This paper presents GraphMend, a high-level compiler that eliminates FX graph breaks in PyTorch 2 programs. Although PyTorch 2 introduced TorchDynamo and TorchInductor to enable just-in-time graph compilation, unresolved dynamic control flow and unsupported Python constructs often fragment models into multiple FX graphs. These fragments force frequent fallbacks to eager mode, incur costly CPU-to-GPU synchronizations, and reduce optimization opportunities. GraphMend addresses this limitation by analyzing and transforming source code before execution. Built on the Jac compilation framework, GraphMend introduces two code transformations that remove graph breaks due to dynamic control flow and Python I/O functions. This design allows PyTorch’s compilation pipeline to capture larger, uninterrupted FX graphs without requiring manual refactoring by developers. Evaluation across eight Hugging Face models shows that GraphMend removes all fixable graph breaks due to dynamic control flow and Python I/O functions, driving the break count to 0 in 6 models and reducing it from 5 to 2 in another model. On NVIDIA RTX 3090 and A40 GPUs, GraphMend achieves up to 75% latency reductions and up to 8% higher end-to-end throughput. These results demonstrate that high-level code transformation is an effective complement to PyTorch’s dynamic JIT compilation pipeline, substantially improving both usability and performance.
I Introduction
@@
PyTorch has emerged as one of the most widely adopted deep learning frameworks in both academia and industry, primarily due to its ease of use and flexibility [1]. However, its general-purpose design incurs efficiency overheads, leaving PyTorch workloads less optimized than handcrafted GPU implementations. To bridge this gap, PyTorch 2 [2] introduced a Python Just-in-Time (JIT) compilation pipeline that translates model operations into optimized GPU code, achieving performance close to handcrafted implementations.
@@
One of the central intermediate representations (IR) in this pipeline is the FX graph [3]. TorchDynamo intercepts and symbolically evaluates Python bytecode to create an FX graph. The FX graph is then passed as the IR to a backend compiler, by default TorchInductor [2], which applies a series of graph-level and kernel-level optimizations before lowering the computation to efficient kernels for GPU execution.
A key component in the PyTorch compilation pipeline is the capture of code into an FX graph through symbolic evaluation. It is important to avoid splitting the forward function into smaller graphs, since unsupported code triggers execution to fall back to eager mode, fragmenting the computation. For example, TorchDynamo often encounters difficulties when resolving conditional jumps, as control flow depending on dynamic values cannot be statically determined [4]. In such cases like dynamic control flow or calls to certain Python built-ins, TorchDynamo cannot capture the affected code into the FX graph. Instead, it inserts a graph break, which splits the computation into multiple fragments [2]. As a result, a single forward function may be compiled into several disjoint FX graphs rather than one unified graph.
@@
The compiled FX graphs run on the optimized backend, while the regions between them execute in standard PyTorch eager mode [1]. Each fallback to eager execution requires GPU kernels to synchronize with the CPU to fetch and schedule subsequent operations. These repeated host-device synchronizations introduce substantial overhead, since control is repeatedly transferred between CPU and GPU instead of executing long fused kernels entirely on the device. Moreover, graph breaks reduce optimization opportunities, as each FX graph is compiled in isolation, preventing cross-graph fusion and global optimization [5, 2].
Avoiding graph breaks is essential for realizing the full performance potential of PyTorch 2. The current PyTorch compilation pipeline begins at the bytecode level and does not exploit source-level information. As a result, when high-level constructs such as conditional branches are lowered into jump instructions at the bytecode level, TorchDynamo often cannot statically resolve them, causing graph breaks. Our key insight is that compilers should not limit transformations to the bytecode level. Instead, more holistic compilation strategies that analyze and transform higher-level program representations closer to the source code can expose additional optimization opportunities. By incorporating source-level analysis and transformation, high-level code regions can be systematically rewritten into forms that enable more effective optimization and unified graph generation within the PyTorch 2 pipeline.
Building on this insight, we introduce GraphMend, a high-level compiler technique for PyTorch 2 models that eliminates two common types of graph breaks those caused by dynamic control flow and Python I/O operations through program analysis and transformation. Unlike standard PyTorch 2, which captures execution at runtime to construct graphs for JIT compilation, GraphMend restructures Python code before execution. We implement GraphMend within the Jac framework [6, 7, 8], a language infrastructure that provides compiler and runtime support for Python programs. The Jac compiler constructs an abstract syntax tree (AST) to capture program structure and a control-flow graph (CFG) to represent execution paths, merging them into a unified IR [9]. Multiple compiler passes operate on this IR to analyze and transform the program, and GraphMend extends this pipeline with additional passes that detect code patterns likely to cause graph breaks. Specifically, GraphMend introduces compiler passes that apply two key transformations: Predicated Dynamic Control Flow and Graph-Epilogue Deferred Side Effects, which systematically eliminate graph breaks. The transformed program is then compiled into standard Python bytecode and executed by the regular CPython interpreter, seamlessly integrating with the existing PyTorch 2 compilation pipeline. This design enables GraphMend to automatically restructure code before it reaches TorchDynamo.
We evaluate GraphMend on a benchmark suite of 8 Hugging Face [10] models that trigger FX graph breaks. GraphMend removes all fixable breaks, reducing the break count to 0 in 6 models and from 5 to 2 in another, with 1 remaining unfixed. On NVIDIA RTX 3090 and A40 GPUs, it achieves 30–75% lower cold-start forward latency, 2.5–25% lower steady-state latency, and 5–8% higher end-to-end throughput. Improvements are larger when a graph break lands in the middle of a hot function where its disruption is highest, and smaller when the break is less disruptive; very small models also tend to see larger relative reductions in per forward pass latency because CPU to GPU handoff dominates a greater share of their runtime.
II Background
Optimization pipelines in PyTorch aim to capture execution graphs that can be globally optimized by compilers, unlike eager mode where each operation runs immediately and independently [1, 3].
II-A Graph Capturing Techniques in PyTorch Before PyTorch 2
Early approaches to graph capture were limited. torch.jit.trace [11] used record–replay at the C++ dispatcher level but ignored Python control flow and often produced incomplete or misleading graphs. torch.jit.script [12] analyzed Python ASTs but followed an “all-or-nothing” rule: if a single construct was unsupported, capture failed entirely. To address these limitations, PyTorch introduced the torch.fx [3] module, enabling symbolic execution of Python code to generate an FX graph, a higher-level IR. Although more flexible, symbolic_trace could still misrepresent certain operations.
II-B FX Graph IR
An FX graph represents operations as nodes connected by data dependencies [2, 3]. Unlike eager execution, which evaluates operations sequentially, an FX graph exposes the computation as a data flow structure. This global view enables the compiler backend to apply optimizations such as operation fusion and kernel specialization.
To illustrate, consider the simple PyTorch forward function in Figure 1(a). When captured by TorchDynamo, it is transformed into the FX graph shown in Figure 1(b), where each operation is represented as a node in the IR.
1def forward(x, y):
2 z = x*2 + y*2
3 return torch.relu(z)
(a) Forward function in PyTorch
%x : torch.Tensor
%y : torch.Tensor
%mul_x : torch.Tensor = torch.ops.aten.mul.Tensor_Scalar(%x, 2)
%mul_y : torch.Tensor = torch.ops.aten.mul.Tensor_Scalar(%y, 2)
%z : torch.Tensor = torch.ops.aten.add.Tensor(%mul_x, %mul_y)
%out : torch.Tensor = torch.ops.aten.relu.default(%z)
return (%out)
(b) Corresponding FX graph IR
Figure 1: A simple forward function (a) and its compiled FX graph IR (b).
This FX graph serves as the foundation for subsequent compiler passes that optimize and lower the computation into efficient GPU kernels.
II-C TorchDynamo in PyTorch 2
PyTorch 2 [2] introduced TorchDynamo, Python bytecode-level tracer tightly integrated with CPython. It symbolically evaluates Python bytecode to extract PyTorch operations into FX graphs, while delegating unsupported constructs back to Python. This avoids incomplete graphs but may still fragment the computation by generating partial graphs by inserting graph-breaks.
@@ II-C TorchDynamo in PyTorch 2
(a) A PyTorch forward pass with data dependent control flow. The conditional branch on x.sum cannot be captured, causing a graph break.
@@
(b) Eliminating the graph break by rewriting data dependent control flow with torch.where, keeping the computation in a single FX graph.
Figure 2: Comparison of control flow handling in torch.compile: (a) graph break due to Python control flow, and (b) fixed version using torch.where.
Figure 3: Profiled traces of forward pass execution across CPU and GPU. (a) Forward pass execution trace of code with graph breaks in Figure 2(a). (b) Forward pass execution trace of equivalent code with graph breaks fixed in Figure 2(b).
III Motivating Example
As discussed in § II, TorchDynamo attempts to capture a single FX graph for the forward function but inserts graph breaks whenever it encounters unsupported constructs, forcing execution to fall back to eager mode [2].
III-A FX graph breaks
Let us consider the modified program in Figure 2(a), which adds a conditional branch to the forward pass of Figure 1(a). When Dynamo tries to symbolically evaluate the function to capture FX graph, it cannot capture the conditional branch because the outcome depends on the runtime value of the tensor expression x.sum. Therefore, it inserts a graph break. As a result, the computation before the branch is recorded in one FX graph, the conditional is executed in eager mode, and the remaining operations are traced into a second FX graph [2, 13].
III-B Impact of graph breaks
The backend compiler (TorchInductor by default) compiles each captured FX graph separately and lowers it to GPU executables such as fused Triton kernels [2, 14]. Each compiled FX graph corresponds to a torch.compiled region. When graph breaks occur, a single Python function results in multiple compiled regions separated by segments of unsupported Python code [13].
@@ III-B Impact of graph breaks
(a) Compiled function with a debug print of tensor statistics. Dynamo inserts a graph break at the print statement.
@@
(b) Fixing the graph break by registering print as a reorderable logging function. TorchDynamo can now safely reorder the call and keep the computation in a single FX graph.
Figure 4: Fixing graph breaks due to Python I/O: (a) direct print causes a graph break, (b) reordering via variable assignment avoids the break.
1 if x > 0:
2 return 1
3 else:
4 return -1
(a) Python function with if-else branch
2 LOAD_CONST 1 (0)
4 COMPARE_OP 4 (>)
6 POP_JUMP_IF_FALSE 14
8 LOAD_CONST 2 (1)
10 RETURN_VALUE
14 LOAD_CONST 3 (-1)
16 RETURN_VALUE
(b) Disassembled Python bytecode
1 if x > 0:
2 return 1
3 return -1
(c) Python function having identical bytecode to (a)
Figure 5: Bytecode-level limitations of TorchDynamo.(a) (a) and (c) show two syntactically different functions with equivalent semantics. (b) Both compile to identical Python bytecode.
During runtime, the first compiled region runs on the GPU, after which control returns to the Python interpreter to execute unsupported code in eager mode before switching back to the GPU for subsequent compiled regions. These context switches introduce overhead and lower GPU utilization per forward pass [2].
This effect is illustrated in Figure 3(a), which shows the CPU–GPU profiling trace of the function in Figure 2(a). The trace reveals two separate torch.compile regions, corresponding to two FX graphs. The graph break is visible in the device-to-host (D2H) memcpy following the first FX graph, where control returns to the CPU to execute unsupported code, leaving the GPU idle until the second CUDA graph launch begins. This fragmentation both increases CPU-side scheduling overhead and prevents cross-graph optimizations such as kernel fusion.
If the graph break is eliminated and the computation is captured in a single FX graph, the execution trace in Figure 3(b) shows one continuous CUDA graph region. In this case, the GPU remains fully utilized without falling back to the CPU, and no device-to-host memory transfers are observed between regions. As a result, kernel launches are streamlined and synchronization overhead is removed. Therefore, eliminating graph breaks is necessary.
IV GraphMend Compiler Technique
@@
As discussed in § III, graph breaks in PyTorch 2 introduce significant overhead, undermining the performance gains expected in PyTorch 2. In this work, we focus on two common sources of graph breaks and introduce a compiler technique that eliminates them prior to execution.
IV-A Classification of Graph Breaks
In practice, most graph breaks arise from two common situations:
  1. 1.
    Data dependent operations [2, 13]: Constructs whose outcome depends on runtime values and hence, cannot be traced.
  2. 2.
    Unsupported operations [2, 13]: Several Python built in functions or certain functionalities that TorchDynamo does not support.
In this section, we discuss these types of graph breaks.
IV-A1 Data dependent operations
Dynamo inserts graph breaks when it encounters control flow that depends on tensor values, such as if-statements or loops, as well as direct tensor data accesses (e.g., .item, .data_ptr). Among these cases, when Dynamo encounters dynamic control flow based on a tensor value as shown in Figure 2(a), we can fix the graph break by rewriting the code into operations supported by the GPU.
IV-A2 Python builtin functions
Some Python built-in functions such as printing, logging, or issuing warnings can also trigger graph breaks. Figure 4(a) shows an example where a debug print statement in the middle of the function causes a break.
IV-B Fixing Common Graph Breaks
Among these graph break reasons, there are two main reasons that are common and fixable which are (1) Data-dependent Control Flow and (2) Python I/O Operations.
IV-B1 Data-dependent Control Flow
Dynamo inserts graph breaks for constructs such as if-statements or loops where branching depend on tensor values. Returning to the example in Figure 2(a), the conditional introduces a graph break. In such cases,we can rewrite the function using tensor operations such as torch.where, which are supported by the GPU. The graph break fixed version using torch.where is shown in Figure 2(b) for the same function in Figure 2(a). This eliminates the graph break and allows Dynamo to capture a single FX graph.
IV-B2 Python I/O Operations
Python I/O functions such as Printing/logging/issuing warnings will result in a graph break. An example for this is shown in Figure 4(a) as there is a debug print statement in the middle of the function. If we can defer print statement to the end of the function, then we can avoid graph breaks as shown in Figure 4(b).
These graph-fixing techniques are simple and straightforward: by applying source-level transformations, we can often avoid graph breaks. However, TorchDynamo traces at the bytecode level, which prevents it from leveraging such transformations and leads to unnecessary graph breaks. For example, consider Figure 5. The function in Figure 5(a) uses an explicit if–else block, while the function in Figure 5(c) omits the else and relies on a fall-through return. Although the two versions differ syntactically, both compile to the same bytecode shown in Figure 5(b). Because TorchDynamo operates purely on bytecode, it cannot distinguish between these semantically equivalent but syntactically distinct patterns.
Similarly, logging operations introduce host I/O into the forward pass. Because it is unsafe to alter such effects directly at the bytecode level, TorchDynamo also treats these sites as graph breaks. In both cases, the limitation stems from relying solely on bytecode-level semantics captured just-in-time.
To address this, we implement a compilation technique that analyzes and transforms programs at a higher-level IR. By operating on high-level IR, we preserve structural information lost in bytecode, enabling more sophisticated analyses and safe source-to-source rewrites that remove graph breaks before execution. This implementation of the technique is shown in Figure 6 which is elaborated in the rest of this section.
IV-C Jac Compilation Pipeline
To implement our compiler technique, we build on the Jac framework [6, 7, 9], which extends Python with additional language features through its runtime system. The Jac compiler accepts both Jac and Python code, parses it into an abstract syntax tree (AST), and progressively lowers it into a unified intermediate representation (UniiR). As illustrated in the top half of Figure 6, UniiR interleaves three structures: AST, control-flow graph (CFG), and symbol table. In this way, structural, semantic, and control-flow information are combined in a single representation. This integration enables cross-cutting analyses that are difficult to achieve when these structures are maintained separately. The CFG is constructed by linking AST nodes to their successors, embedding control-flow edges within the tree. This process is implemented in the CFGBuildPass.
Our GraphMend pass takes UniiR as input and performs analyses to identify the two types of graph breaks, as illustrated in the bottom half of Figure 6. It then applies transformations on the AST to rewrite the code and eliminate these breaks. Afterward, UniiR is rebuilt and verified before being passed back to the Jac compiler pipeline. At that stage, the pipeline can either compile the program to bytecode for execution on the Python interpreter or generate a transformed Python source file.

### v2

@@ Abstract
This paper presents GraphMend, a high-level compiler technique that eliminates FX graph breaks in PyTorch 2 programs. Although PyTorch 2 introduced TorchDynamo and TorchInductor to enable just-in-time graph compilation, unresolved dynamic control flow and unsupported Python constructs often fragment models into multiple FX graphs. These fragments force frequent fallbacks to eager mode, introduce costly CPU-to-GPU synchronizations, and reduce optimization opportunities. GraphMend addresses this limitation by analyzing and transforming source code before execution. Built on the Jaseci compilation framework, GraphMend introduces two code transformations that remove graph breaks due to dynamic control flow and Python side effects. This design allows PyTorch’s compilation pipeline to capture larger, uninterrupted FX graphs without requiring manual refactoring by developers. Evaluation across eight Hugging Face models shows that GraphMend removes graph breaks due to dynamic control flow and Python side effects, reducing the break count to 0 in 6 models and reducing it from 5 to 2 in another model. On NVIDIA RTX 3090 and A40 GPUs, GraphMend achieves up to 75% latency reductions and up to 8% higher end-to-end throughput. These results demonstrate that high-level code transformation is an effective complement to PyTorch’s dynamic JIT compilation pipeline, substantially improving both usability and performance.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
@@
PyTorch has emerged as one of the most widely adopted deep learning frameworks in both academia and industry, primarily due to its ease of use and flexibility Paszke et al. (2019). However, its general-purpose design incurs efficiency overheads, leaving PyTorch workloads less optimized than handcrafted GPU implementations. To bridge this gap, PyTorch 2 Ansel et al. (2024) introduced a Python Just-in-Time (JIT) compilation pipeline that translates model operations into optimized GPU code, achieving performance close to handcrafted implementations.
@@
One of the central intermediate representations (IR) in this pipeline is the FX graph Reed et al. (2022). TorchDynamo intercepts and symbolically evaluates Python bytecode to create an FX graph. The FX graph is then passed to a backend compiler, by default TorchInductor Ansel et al. (2024), which applies a series of graph-level and kernel-level optimizations before lowering the computation to efficient kernels for GPU execution.
When tracing the FX graph, TorchDynamo often struggles with unsupported Python operations, which it handles by falling back to eager mode. These unsupported operations include dynamic control flow such as conditionals that depend on runtime values and certain Python side effects like I/O operations. In such cases, it inserts graph breaks, resulting in multiple disjoint FX graphs instead of a single unified one Zheng et al. (2023); Ansel et al. (2024).
@@
These disjoint FX graphs run on the optimized backend, while the regions between them execute in standard PyTorch eager mode Paszke et al. (2019). Each fallback to eager mode forces GPU kernels to synchronize with the CPU to schedule subsequent operations, introducing significant overhead from repeated host-device communication. These synchronizations not only fragment execution but also reduce optimization opportunities, as each FX graph is compiled independently, preventing cross-graph fusion and global optimization Agrawal et al. (2019); Ansel et al. (2024).
Avoiding graph breaks is essential for realizing the full performance potential of PyTorch 2. Some of these breaks can be eliminated by transforming Python code into forms that use PyTorch-supported operations, allowing the compiler to capture more of the program as a single, continuous FX graph. However, the current compilation pipeline analyzes code at the bytecode level, operating instruction by instruction without understanding higher-level program structures such as control flow, data dependencies, and side effects. Our key insight is that analyzing and transforming higher-level program representations closer to the source code, and integrating this with bytecode-level analysis, enables a more holistic compilation strategy that exposes richer optimization opportunities.
@@
Figure 1: A PyTorch forward pass with data-dependent control flow. The conditional branch on x.sum cannot be captured, causing a graph break. (See also Figure 2 for the transformed version that eliminates the break.)
Building on this insight, we introduce GraphMend, a high level compiler technique for PyTorch 2 that eliminates two major types of graph breaks, those arising from dynamic control flow and several Python side effects, through high-level program analysis and transformation. GraphMend extends the standard PyTorch 2 compilation pipeline with an AST-level transformation phase that restructures Python code before bytecode generation, improving traceability by eliminating graph breaks.
We implement GraphMend within the Jaseci framework, built on Jac language Mars et al. (2023); Mars (2025); Jaseci Labs (2025), which is a super-set of Python. The Jac compiler constructs an abstract syntax tree (AST) and a control flow graph (CFG), merging them into a unified intermediate representation (UniiR) Dantanarayana et al. (2025). Multiple compiler passes analyze and transform this IR, and GraphMend extends this pipeline with additional passes that detect and restructure code patterns prone to graph breaks. Specifically, it introduces two transformations, Predicated Dynamic Control Flow and Graph Epilogue Deferred Side Effects, which systematically eliminate these breaks.
We evaluate GraphMend on a benchmark suite of 8 Hugging Face Hugging Face (2025) models that trigger FX graph breaks. GraphMend removes all fixable breaks, reducing the break count to 0 in 6 models and from 5 to 2 in another, with 1 remaining unfixed. On NVIDIA RTX 3090 and A40 GPUs, it achieves 30–75% lower cold-start forward latency, 2.5–25% lower steady-state latency, and 5–8% higher end-to-end throughput.
@@
Figure 2: Eliminating the graph break by rewriting data-dependent control flow with torch.where, keeping the computation in a single FX graph. Together with Figure 1, these illustrate how dataflow rewriting replaces dynamic Python branching with predicated tensor execution.
Figure 3: Profiled traces of forward pass execution across CPU and GPU. (a) Forward pass execution trace of code with graph breaks in Figure 1. (b) Forward pass execution trace of equivalent code with graph breaks fixed in Figure 2.
(a) Existing graph-break in Phi-4-mini instruct model built by Microsoft.
(b) Version of Phi-4-mini instruct model with the graph break fixed, without affecting program correctness.
Figure 4: (a) Existing graph-breaks in a real model and (b) how it can be fixed to produce a single contiguous FX graph.
2 Motivation
FX graph capturing:
TorchDynamo Ansel et al. (2024) captures graphs by symbolically interpreting Python bytecode, extracting PyTorch operations into FX graphs while falling back to eager mode for unsupported code.
FX graph breaks.
Let us consider the program in Figure 1, which has a conditional branch. When Dynamo tries to symbolically evaluate the function to capture FX graph, it cannot capture the conditional branch because the outcome depends on the runtime value of the tensor expression x.sum. Therefore, it inserts a graph break. As a result, the computation before the branch is recorded in one FX graph, the conditional is executed in eager mode, and the remaining operations are traced into a second FX graph Ansel et al. (2024); PyTorch Contributors (2025).
Impact of graph breaks.
When graph breaks occur, the function is divided into multiple FX graphs, each compiled separately into GPU kernels Ansel et al. (2024); Ghosh et al. (2025). Each FX graph forms a distinct torch.compiled region separated by segments of unsupported Python code PyTorch Contributors (2025).
During execution, GPU runs the first compiled region. Control then returns to the Python interpreter to execute the unsupported code in eager mode and switches back to GPU for the next compiled region Ansel et al. (2024). These context switches introduce CPU–GPU synchronization overhead and reduce GPU utilization.
Figure 3(a) illustrates this effect for the function in Figure 1. The trace shows two separate torch.compile regions divided by a device-to-host (D2H) memory transfer, where the GPU remains idle until the next CUDA graph launch. This fragmentation increases CPU scheduling overhead and prevents optimizations such as kernel fusion. When the break is removed and captured as a single FX graph, the execution trace in Figure 3(b) shows one continuous CUDA region with full GPU utilization and no intermediate D2H transfers. Kernel launches become streamlined, and synchronization overhead is eliminated.
Real world model example with graph breaks.
Phi-4-mini Instruct Microsoft (2025) is a lightweight open-source instruction-following model from Microsoft, available on Hugging Face as an efficient alternative to larger transformers. According to Hugging Face statistics Hugging Face (2025), the model recorded over 221k downloads in September 2025, reflecting strong community adoption.
To examine the impact of graph breaks in a widely used model, we profiled Phi-4-mini Instruct using PyTorch 2 and identified 5 graph breaks caused by data-dependent control flow. Figure 4(a) illustrates this case, where conditional logic depending on runtime tensor values introduces a graph break.
We rewrote the same control flow using PyTorch-compatible transformations with the torch.where operation to evaluate both branches within the FX graph while correctly handling side effects, as shown in Figure 4(b). After applying this modification, all graph breaks were eliminated. Profiling showed over a 50% latency reduction per forward pass in cold runs across multiple GPUs and more than a 5% improvement in steady-state runs compared to the version with graph breaks (further discussed in § 5), by reducing the impact of the graph breaks. This example demonstrates that targeted code transformations can effectively remove graph breaks and substantially improve performance.
3 Understanding and Eliminating Common Graph Breaks
@@
Figure 5: Compiled function with a print statement. Dynamo inserts a graph break at the print statement. (See also Figure 6 for the transformed version that avoids the break.)
@@
Figure 6: Fixing the graph break in Figure 5 by registering print as a reorderable logging function. The value to be printed is assigned to a buffer during tracing and flushed at the end.
@@
In this work, we focus on two common causes of graph breaks, outlining code transformations to eliminate them and explaining why bytecode-level fixes are challenging without higher-level IR context.
Data dependent operations.
Ansel et al. (2024); PyTorch Contributors (2025) In PyTorch 2, symbolic tracing builds FX graphs using FakeTensor objects, assuming control flow decisions can be inferred from static information such as tensor shapes or constants. When a condition depends on actual tensor data, (for example, if x.sum() > 0: in Figure 1), the compiler cannot safely evaluate it, so PyTorch inserts a graph break.
We can fix them by expressing data-dependent branches using tensor operations such as torch.where or torch.cond, which are supported on GPU. Figure 1-2 shows the rewritten version of the same function, where the use of torch.where eliminates the graph break.
Python Side-effects.
Side effecting operations such as print or logging calls also cause graph breaks because they interact with the Python runtime and cannot be represented in a functional graph. Deferring these effects until after the graph executes can remove the break while preserving program behavior. In Figure 5 and 6, we show how moving a print statement to the end of the function allows the entire function to remain in one FX graph.
Why automating these transformations at the bytecode level is challenging?
Automating these transformations must occur before PyTorch’s bytecode analysis, but performing safe transformations at the bytecode level is challenging since much of the high-level semantic information is lost during compilation.
When high-level constructs such as if/else or while loops are lowered to bytecode, they become low-level jump instructions like POP_JUMP_IF_FALSE and JUMP_BACKWARD. Since these opcodes are reused across different control-flow patterns, distinguishing loops from conditionals and identifying their bodies is non-trivial and often requires partial reverse engineering of the bytecode. Ensuring correctness during transformation further complicates this process.
Moreover, performing data-flow analysis at the bytecode level is difficult since Python uses a stack-based rather than a register-based model. While variable operations themselves are explicit, intermediate expression values are pushed and popped from the operand stack without explicit naming, obscuring the flow of computed values without symbolic stack simulation. This simulation adds complexity and overhead to the analysis, whereas analyses on higher-level AST representations with explicit variable bindings and control flow are far simpler.
Handling side effects at the bytecode level is also challenging. While delaying operations like built-in print or logger calls may be safe, Python’s dynamic nature allows users to redefine these functions. Determining whether a call refers to the original built-in or a user-defined override requires tracking symbol bindings, which is only possible with higher-level semantic context.
A higher-level IR that retains program structure and data dependencies is therefore better suited for automatically applying these transformations.
4 GraphMend Compiler Technique
In this section we introduce our compiler technique GraphMend, which automates these code transformations using AST and CFG level analyses and AST-transformations, to eliminate graph breaks. The implementation of this technique is shown in Figure 7 which is elaborated in the rest of this section.
4.1 Jac Compilation Pipeline
To implement our compiler technique, we build on the Jaseci framework Mars et al. (2023); Mars (2025); Dantanarayana et al. (2025), a compiler stack designed for advanced analysis and transformation of Python programs. Jaseci compiler accepts both Jac and Python code, parses them into an AST, and progressively lowers the representation into UniiR. As shown in the top half of Figure 7, UniiR unifies the AST, CFG and symbol table into a single structure, combining syntactic, semantic, and control-flow information.